The resistance of an exercise bike is often provided by a generator; that is, the energy that you expend is used to generate electric energy, which is then dissipated. Rather than dissipate the energy, it could be used for practical purposes.

A) A typical person can maintain a steady energy expenditure of 400 W on a bicycle. Assuming a typical efficiency for the body and a gerator that is 85% efficient, what useful electric power could you proudce with a bicycle-powered generator?

P=???

B) How many people would need to ride bicycle generators simultaneously to power a 340 W TV in the gym?

N=???
Answer #1

A)
400 * 0.25 * 0.85 = 85 W

B)
340 / 85 = 4 persons
Answer #2

1)P=85*400/100=340W

2)1 person is enough as seen above
Answer #3

A)
400 * 0.25 * 0.85 = 85 W

B)
340 / 85 = 4 persons
Answer #4

Part A

(400 W) (85%) = 340 W

Part B

(340W)/(340 W) = 1 people